OpenStudy (anonymous):

A student at harvard University decided to earn money by using his car as a taxi. It costs the student $693.00 to insure his car for the 4 months he drove as a taxi driver. He spends $452 per month on gas. If he lives wit his parents and has no other expenses for those 4 months and charges an average of $7.00 per fare, How many fares will he have to get to be able to pay his expenses and his tuition of 3280?

OpenStudy (anonymous):

469 fares 835.86 fares 528.85 fares None of the above

OpenStudy (anonymous):

I'm just checking my work

OpenStudy (anonymous):

I think it's none of the above

OpenStudy (thesecret20111):

His expenses for the four months = 693+452*4+3280 = 5781

OpenStudy (anonymous):

So b?

OpenStudy (thesecret20111):

He earns $7 a fare so 5781/7 = amount of fares needed

OpenStudy (anonymous):

Yeah it's none of the above

OpenStudy (thesecret20111):

I'm not sure, that would be $825.85 I think its none of the above

OpenStudy (thesecret20111):

I hope I haven't made a small error somewhere

OpenStudy (thesecret20111):

I didn't mean $ above, I meant number of fares

OpenStudy (anonymous):

I miscalculated and I just needed to make sure

OpenStudy (anonymous):

Thanks

OpenStudy (thesecret20111):

Ok np :)

OpenStudy (anonymous):

b was pretty close which is why I thought it b :P

OpenStudy (thesecret20111):

Yeah, b has me unsure whether or not I've made a small error :/

OpenStudy (anonymous):

No you haven't, my teacher flutters with me like that all the time

OpenStudy (thesecret20111):

Ok :P

OpenStudy (thesecret20111):

Happy to help!